Explain the impedence in LCR circuit.


1
Expert's answer
2021-01-28T12:19:11-0500

The impedance of LCR cirquit in Cartesian form is:


"Z=R+j\omega L+\frac {1}{j\omega C}=R+j(\omega L-\frac {1}{\omega C})."


The real part of impedance (R) is called resistive impedance, the imaginery part is reactive impedance, which consists of inductive ("\omega L") and capasitive ("\frac{1}{\omega C}") components.


The impedance of LCR cirquit can also be expressed in the polar form:


"Z=|Z|e^{jarg(Z)}" ,


where |Z| - the magnitude of impedanse, which represents the ratio of the voltage difference amplitude to the current amplitude and equals:


"|Z|=\sqrt{R^2+(\omega L-\frac{1}{\omega C})^2}"


arg(Z) - the argument of impedanse, which represents the  phase difference between voltage and current and equals:


"arg(Z)=arctan\frac{\omega L-\frac{1}{\omega C}}{R}."